OpenStudy (anonymous):

A ladder 15 feet long leans against a wall and forms an angle of 45o with the ground. How far from the wall is the ladder? helppp please!!!

OpenStudy (anonymous):

cos 45 = x/ 15 solve fo x

OpenStudy (akshay_budhkar):

u can do that right??

myininaya (myininaya):

Why you post this twice?

OpenStudy (anonymous):

 \[\cos 45=Adjacent / Hypotenuse\]\[Adjacent = 15 * \cos45 = 15 * 0.7071 = 10.606\] get free math homework help at http://www.mathskey.com